In order to complete your case analysis successfully, you must

  • identify the role you are playing,
  • assess user needs
  • analyze user needs or issues (qualitatively and quantitatively), and
  • provide a recommendation and conclusion.

An average grade will result from answering all questions with basic coverage and accuracy, showing all your work. Additional points come from including greater detail, astute, informed commentary where appropriate and connections to readings and other content.

Case: Cost Structures for Global Shippers Inc.

Management from Global Shippers Inc, an international shipping business, is in the process of assessing the choice between two different cost structures for the business. Option A has relatively higher variable costs per unit shipped but lower annual fixed costs, while Option B has the opposite—relatively lower variable costs in its cost structure but higher fixed costs. Assume that delivery selling prices per unit are constant. The table below contains critical information in making the decision:

Cost Information

Option A

Option B

Delivery price (revenue) per shipment

$100

$100

Variable cost per shipment delivered

$85

$60

Contribution Margin per unit

$15

$40

Fixed costs (annual)

$1,200,000

$4,500,000


Management wants you to write a professional report, answering the following questions:

Questions

1) What is the break-even point, in terms of volume (i.e., number of shipments per year), for Option A? Option B?

(2) How many shipments would have to be made under Option A to produce operating income of $30,000 for an annual period?

(3) How many shipments per year would have to be made under Option A to produce an operating margin equal to 9% of sales revenue?  

(4) How many shipments are required under Option B to produce net income of $180,000 per year, given a corporate tax rate of 40%?

(5) Assume that for the coming year total fixed costs are expected to increase by 15% for each of the two options. What is the new break-even point, in terms of number of shipments, for each option? By what percentage did the break-even point change for each case? How do these figures compare to the percentage increase in budgeted fixed costs?

(6) Assume an average income-tax rate of 20%. What volume (number of shipments) would be needed to generate net income of 5% of revenue for each option?  

(7) Which option do you think is the more profitable one for this business? Explain.

(8) Which option do you consider to be more risky to the business? Explain (calculate degree of operating leverage to help answer this question).
Answer #1

Solution

Global Shippers Inc

  1. Break-even point in terms of volume for

Option A –

Break-even point, in terms of volume = fixed cost/unit contribution margin

Fixed cost = $1,200,000

Contribution margin per unit = $15

Break-even points in terms of volume = 1,200,000/15 = 80,000 shipments

Option B –

Break-even point, in terms of volume = fixed cost/unit contribution margin

Fixed cost = $4,500,000

Contribution margin per unit = $40

Break-even points in terms of volume = 4,500,000/40 = 112,500 shipments

  1. Number of shipments needed –

Option A – to produce operating income of $30,000 for an annual period:

Desired number of shipments = (target income + fixed cost)/unit contribution margin

Target income = $30,000

Fixed cost = $1,200,000

Unit contribution margin = $15

Desired number of shipments = (30,000 + 1,200,000)/15 = 82,000 shipments

  1. Number of shipments needed to earn operating margin equal to 9% of sales revenue:

Option A – to produce operating margin equal to 9% of sales revenue.

Desired number of shipments = (target income + fixed cost)/unit contribution margin

Assuming the desired number of shipments to be A,

Target income = 9% x $100 x A = 9A

Fixed cost = $1,200,000

Unit contribution margin = $15

Desired number of shipments = (9A + 1,200,000)/15 = A shipments

15A = 9A + 1,200,000

6A = 1,200,000

A = 1,200,000/6 = 200,000 units

Target income = 9% x $100 x 200,000 = $1,800,000

  1. Number of shipments required under Option B to produce net income of $180,000 per year, given a corporate tax rate of 40%:

Net income =180,000

Tax rate = 40%

Before tax income = 180,000/60% = $300,000

Fixed cost = $4,500,000

Contribution margin per unit = $40

Number of shipments = (300,000 +4,500,000)/40 = 120,000 shipments

Hence, shipments needed to produce net income of $180,000 per year under option B = 120,000.

  1. Determination of new break-even point when fixed cost increases by 15% for both the options:

Option A –

Fixed cost = $1,200,000

Add increase = 15% x 1,200,000 = 180,000

New fixed cost = 1,380,000

Contribution margin per unit = $15

Break-even point, in terms of number of shipments = 1,380,000/15 = 92,000 shipments

Change in break-even point –

Original break-even point 80,000 shipments

New break-even point in terms of volume = 92,000

Increase = 92,000 – 80,000 = 12,000 shipments

Percent increase in break-even point, volume = 12,000/80,000 = 15%

Option B –

Fixed cost = $4,500,000

Add increase = 15% x 4,500,000 = 675,000

New fixed cost = 5,175,000

Contribution margin per unit = $40

Break-even point, in terms of number of shipments = 5,175,000/40 = 129,375 shipments

Original beak-even point in volume = 112,500 shipments

New break-even point , volume = 129,375 shipments

Increase = 129,375 – 112,500 = 16,875 shipments

Percent increase = 16,875/112,500 = 15%

The increase in break-even point, in terms of shipment for both the options is 15%, which is equal to the percent increase in fixed cost for both the options. This indicates that the break-even point, in terms of value increases in proportionately with the fixed cost.

  1. Shipments needed to generate net income of 5% revenue for each option, assuming a tax rate of 20%:

Option A -

Net income = 5% x $100 x A = $5A

Tax rate = 20%

Before tax income = 5A/80% = $6.25A

Fixed cost = $1,200,000

Contribution margin per unit = $15

Number of shipments = (6.25A +1,200,000)/15 = A

15A = 6.25A +1,200,000

8.75 A = 1,200,000

A = 1,200,000/8.75 = 137,143 shipments

Option B –

Net income = 5% x $100 x A = $5A

Tax rate = 20%

Before tax income = 5A/80% = $6.25A

Fixed cost = $4,500,000

Contribution margin per unit = $40

Number of shipments = (6.25A +4,500,000)/40 = A

40A = 6.25A +4,500,000

33.75 A = 4,500,000

A = 4,500,000/33.75 = 133,333 shipments

Hence, shipments needed to produce net income of 5% of sales revenue per year under option B = 133,333.

  1. The most profitable options:

Option A is the most profitable option for this business. The break-even point, in terms of volume for option A (80,000) is less compared to the break-even point, in terms of volume for option B (112,500). The lesser the break-even point, the quicker the company would start making profits.

  1. More risky option –

The option B is probably the more risky option, in view of higher proportion of fixed costs.